OpenStudy (anonymous):

Consider the function y(θ) = 10 + 4cos(9θ) What is the largest value of the function?

OpenStudy (anonymous):

think about what the cos function does

OpenStudy (anonymous):

Since \[-1\le \cos \theta \le1\] Largest value is when \[\cos 9\theta = 1\]